I need help with this problem
The first part of a trip a canoeist travels 59 miles. the canoeist travels 17 miles on the second part of the trip at a speed 5mph slower. The total time for the trip is 4hrs. What was the speed on each part of the trip?
Help!
Thanks
Gina

Answer by richwmiller(17219)

r*t=d
r*x=59
(r-5)*y=17
x+y=4
three equations three unknowns
discard negative answers
r = 1/2 (24+sqrt(281)), x = -2/5 (sqrt(281)-24), y = 2/5 (sqrt(281)-14)
x=2.89478, y=1.10522
r=20.3815 first part
r=15.3815 second part